OpenStudy (anonymous):

if each of four numbers is trippled, their average is multiplied by a) 12 b) 3 c) 4 d) 9 e) 81

OpenStudy (kinggeorge):

Take an example. Suppose all four numbers were 1. Then the average would be 1. If they were all tripled, then they would all be 3, and the average would be 3 as well.

OpenStudy (kinggeorge):

So you also have to triple the average.

OpenStudy (anonymous):

well how do i calucate is : 3 * 3 * 3 = 27 4 * 4 * 4 = 64 9 * 9 * 9 = 729 all final values are divisbile by 3 does it make sense ... to solve it like this

OpenStudy (kinggeorge):

If you would like an algebraic explanation, suppose \(a, b, c, d\) are your 4 numbers. Then the average is \[{a+b+c+d \over 4}\]Now we triple the numbers and find the average again.\[{3a+3b+3c+3d \over 4} ={3(a+b+c+d) \over 4}=3\cdot{a+b+c+d \over 4}\]

OpenStudy (anonymous):

Excellent ... Thanks got it...

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

How can I post another question ?

OpenStudy (kinggeorge):

Click the button that says "Close," and you'll be able to type another one.